OpenStudy (anonymous):

Express the interval (−2, 6) in terms of an inequality involving absolute value. (Use the variable x to describe the interval.)

OpenStudy (tkhunny):

Find the center of the interval.

OpenStudy (anonymous):

i have no idea how to do this

OpenStudy (tkhunny):

Did you do what I said?

OpenStudy (anonymous):

whats the center 4?

OpenStudy (tkhunny):

Why would that be the center? It's not +2, it's -2. You MUST pay better attention.

OpenStudy (anonymous):

-2+6=4?

OpenStudy (anonymous):

-2-6 = -8

OpenStudy (tkhunny):

That's the sum. That's not the average or center of two values.

OpenStudy (tkhunny):

Find the center of the interval. You can just walk ir off if you like. -2 6 One step each way -1 5 0 4 1 3 2 2 <== There it is!

OpenStudy (debbieg):



OpenStudy (anonymous):

how will i express that with absolute value?

OpenStudy (tkhunny):

How far is the center from the two endpoints?

OpenStudy (anonymous):

(0,4)?

OpenStudy (tkhunny):

Why would it be an interval? We need an absolute distance. How far is 2 from 6? How far is 2 from -2?

OpenStudy (anonymous):

4 and 2

OpenStudy (tkhunny):

The difficulty with what I am dragging you through is that we are BUILDING the solution. You cannot see the end from where we are. I hope to show you a PROCESS, not an answer. Where is the center of the interval? (6 - (-2))/2 = 2 How far is each endpoint from the center? 6-2 = 4 and -2 - (-2) = 4 This is how absolute values work. It is sometimes called "magnitude" or "absolute distance" or "norm" and maybe a few other things. In any case, it is a positive number referring to how far apart things are. Our task here is the describe the set of all numbers that are less than 4 from 2. How do we do that?

OpenStudy (anonymous):

2<x<4?

OpenStudy (tkhunny):

You're just guessing and you failed to use an absolute value. 4 is the ONLY distance that matters. 2 is the ONLY center of the interval that matters. Give it another go. Hint: It will have only one inequality.

OpenStudy (anonymous):

|x|+2<4

OpenStudy (tkhunny):

Now THAT was a good effort. |x-2| < 4 Read this, everything that is less than 4 away from 2.

OpenStudy (anonymous):

Thanks alot for the help.. I really appreciate it!